OpenStudy (anonymous):

can someone plx give me the answer. The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? x = 1, x = 17 x = −7, x = 9 x = 7 x = 9 No solution

OpenStudy (anonymous):

you can manipulate this equation the same way you can ANY equation, until you deal with the absolute value. Go ahead and do that.

OpenStudy (anonymous):

"That" meaning, isolate the |x-8|

OpenStudy (campbell_st):

subtract 15 from both sides of the equation 3| x - 8| = 3 divide both sides by 3 and you are now solving 2 equations x - 8 = 1 and x - 8 = -1 hope it helps

OpenStudy (anonymous):

so would it be the second one

OpenStudy (campbell_st):

well one answer is the max temp and 1 is the min temp.. you need 2 answers

OpenStudy (anonymous):

so no soultion

OpenStudy (campbell_st):

well you could try to solve the 2 equations x - 8 = 1 and x - 8 = -1 simply add 8 to both sides of the equation...

OpenStudy (anonymous):

I tyed and did not get it

OpenStudy (campbell_st):

ok... well if you follow the instruction of adding 8 to both sides of each equation you'll get the solution

OpenStudy (anonymous):

3erd